# EVTAR: End-to-End Try on with Additional Unpaired Visual Reference

### v1

@@ 4.2 QUALITATIVE COMPARISON
Although the previous method can fit the garment onto a human body, the generated results often fail to accurately capture the garment’s cut, style, details, and other characteristics that determine how it should be worn. In contrast, our method achieves superior visual fidelity compared to baselines. As shown in Fig. 6, on the VITON dataset, our EVTAR produces a highly realistic rendering of challenging garment materials such as hollow or semi-transparent fabrics without relying on reference images. For example, in the case of garments with hollow lace structures, our model accurately preserves the delicate perforated patterns and material transparency. In contrast, the baselines fail to reconstruct these details, often generating incorrect solid textures or spurious dotted patterns. Moreover, our approach demonstrates the best preservation of garment patterns: the printed letters and logos on clothes remain clear and consistent with the input, highlighting its advantage in maintaining fine-grained appearance details. The generation quality is further enhanced with the addition of reference images, which intuitively verify the effectiveness of the proposed tryon method. Even without using agnostic masks and by directly transferring the clothing $\bm{c}_{i}$ onto person images $\bm{\bar{p}}_{i,\bm{c}_{j}}$, our method still outperforms most baseline approaches.
@@
4.3 Training on VRF Dataset and Evaluation
Table 4: We compare our approach with [8] under cross-dataset evaluation, reporting results on both paired and unpaired settings for the VITON-HD and DressCode datasets. The best results are highlighted in bold.The * marker refers to the results reported in previous work.The symbol “+R” denotes results with reference image input, and “MF” indicates mask-free input images. Subscripts $p$ and $u$ represent the paired and unpaired test settings, respectively.
@@ 4.3 Training on VRF Dataset and Evaluation
To better evaluate the effectiveness of our proposed end-to-end virtual try-on framework and data construction method, we construct a mixed dataset, Mixed-Virtual-Ref, by combining data from DressCode, VITON-HD, FashionTryOn [55], ViViD [27], and IGPairs [12]. We carefully select high-quality samples using Qwen2.5-VL to filter out low-quality target images, generate cloth-agnostic images with the mask generator from [14], and produce reference and unpaired person images using our data engine. We collect 103,936 image pairs to train our end-to-end virtual try-on model. The training hyperparameters are consistent with those used in the DressCode and VITON-HD experiments. We also evaluate the model on the DressCode and VITON-HD test sets, with quantitative results reported in Table 4. Our method (EVTAR) consistently achieves superior or comparable performance to OOTDiffusion [8] across both VITON-HD and DressCode benchmarks under paired and unpaired settings. Although EVTAR is not explicitly trained on DressCode or VITON-HD, the model trained on the mixed VRF dataset outperforms dataset-specific baselines across most metrics (e.g., FID and LPIPS), demonstrating a strong generalization capability. These results highlight the effectiveness of our approach in learning transferable representations that maintain robustness across diverse datasets.
@@ 5 Conclusion
In this work, inspired by the human painting process, we design a diffusion model that generates high-quality, class-conditional samples without relying on time or class embeddings by mitigating manifold overlap through geometric decoupling of timesteps. The proposed approach simplifies diffusion models and effectively overcomes the challenges introduced by removing time conditioning, while preserving generation performance. More broadly, we offer a novel geometric perspective to understand the dynamics of diffusion processes in high-dimensional data spaces, inspired by the inherent sparsity of data distributions in high-dimensional statistics. We anticipate that our theoretical insights will inspire future research to simplify diffusion-based models by eliminating time and class embeddings and reducing their training and parametric complexity.

### v2

@@ 4.2 QUALITATIVE COMPARISON
Although the previous method can fit the garment onto a human body, the generated results often fail to accurately capture the garment’s cut, style, details, and other characteristics that determine how it should be worn. In contrast, our method achieves superior visual fidelity compared to baselines. As shown in Fig. 6, on the VITON dataset, our RefVTON produces a highly realistic rendering of challenging garment materials such as hollow or semi-transparent fabrics without relying on reference images. For example, in the case of garments with hollow lace structures, our model accurately preserves the delicate perforated patterns and material transparency. In contrast, the baselines fail to reconstruct these details, often generating incorrect solid textures or spurious dotted patterns. Moreover, our approach demonstrates the best preservation of garment patterns: the printed letters and logos on clothes remain clear and consistent with the input, highlighting its advantage in maintaining fine-grained appearance details. The generation quality is further enhanced with the addition of reference images, which intuitively verify the effectiveness of the proposed tryon method. Even without using agnostic masks and by directly transferring the clothing $\bm{c}_{i}$ onto person images $\bm{\bar{p}}_{i,\bm{c}_{j}}$, our method still outperforms most baseline approaches.
@@
Table 3: We compare our approach with [xu2025ootdiffusion] under cross-dataset evaluation, reporting results on both paired and unpaired settings for the VITON-HD and DressCode datasets. The best results are highlighted in bold. The * marker refers to the results reported in previous work. The symbol “+R” denotes results with reference image input, and “MF” indicates mask-free input images. Subscripts $p$ and $u$ represent the paired and unpaired test settings, respectively.
@@
4.3 Training on VRF Dataset and Evaluation
To better evaluate the effectiveness of our proposed person-to-person virtual try-on framework and reference data construction pipeline, we train our RefVTON model on the VRF dataset collected using our data engine introduced in sec 3.4. The training hyperparameters are consistent with those used in the DressCode and VITON-HD experiments. For evaluation, we quantitatively test the generated try-on images on the DressCode and VITON-HD test sets, with quantitative metrics reported in Table 3. Our method (RefVTON) consistently achieves superior or comparable performance to OOTDiffusion [xu2025ootdiffusion] across both VITON-HD and DressCode benchmarks, in both paired and unpaired settings. Although RefVTON is not explicitly trained on DressCode or VITON-HD, the model trained on the mixed VRF dataset outperforms dataset-specific baselines across most metrics (e.g., FID and LPIPS), demonstrating a strong generalization capability. These results highlight the effectiveness of our approach in learning transferable representations that maintain robustness across diverse datasets.
Figure 8: Qualitative results of the ablation study across different settings. “Ref.” denotes that a reference image is provided, while “MF” indicates mask-free inputs using the original person image instead of a masked agnostic image.
4.4 Ablation Study
We conduct an ablation study to examine our model under four settings (w/&w/o mask, w/&w/o Ref.). As shown in Table 1, our model maintains consistently strong performance across all settings. Introducing a reference image yields clear improvements in both mask-based and mask-free modes, while moving from mask-based to mask-free inputs causes only mild metric fluctuations, confirming the model’s stable robustness without masks.
Fig. 8 further provides qualitative comparisons. The first two rows show that reference images help the model correctly infer garment structures and materials that are ambiguous in the flat images (e.g., hollow textures, semi-transparency). The last two rows illustrate that mask quality heavily affects mask-dependent models: overly aggressive masks remove important items (e.g., handbags), while conservative masks retain unwanted regions (e.g., legs), leading to incorrect garment geometry. In contrast, our mask-free model consistently produces correct outputs regardless of mask or reference conditions, demonstrating that mask-free capability reduces reliance on mask quality and enables more flexible and stable try-on performance.
@@ 5 Conclusion
This paper introduces RefTON, a virtual try-on framework that supports both mask-based and mask-free inference and leverages reference images to guide the try-on process. We extend Flux-Kontext to handle multi-condition inputs of varying resolutions via a modified positional index. To train RefTON, we propose a reference data generation pipeline integrating Qwen2.5-VL and Flux-Kontext. This design allows RefTON to faithfully preserve translucent fabrics, intricate designs, and fine details, consistently outperforming existing methods both quantitatively and qualitatively, achieving state-of-the-art performance in all settings.